OpenStudy (anonymous):

pyramid has a volume of 24 and a base area of 12. What is the length of the altitude of the pyramid?

OpenStudy (anonymous):

v=1/3AT T=6

OpenStudy (anonymous):

(1/3)*b*h = 24 (1/3)*12*h = 24 4*h = 24 h = 24/4 h = 6